Two particles, each of mass 0.000290kg and speed 5.46 m/s, travel in opposite directions along parallel lines separated by 4.20 cm. (a) What is the magnitude L of the angular momentum
of the two-particle system around a point midway between the two
lines? (b) Does the value of L change if the point about which it is
calculated is not midway between the lines? If the direction of
travel for one of the particles is reversed,what would be (c) the an-
swer to part (a) and (d) the answer to part (b)?
1
Expert's answer
2012-08-14T11:25:42-0400
a) angular momentum is L=m1*v1*r1+m2*v2*r2=0.000290*5.46*0.021 + 0.000290*5.46* 0.021 =0.0000665028 kg*m^2/s.
b) yes, it would change because the two values, calculated in the a) part would be different (different r1 and r2).
c) Then answer changes because in this case we subtract two equal values& L=m1*v1*r1-m2*v2*r2=0.000290*5.46*0.021 - 0.000290*5.46* 0.021 =0.
d) In this case we also have subtraction, but not 0 in result because the two values are different (r1 and r2 different).